Section 4-Seare Resourees20 Kaiser Company produces two products and selected data is shown Products 2 s 50 s 15 2,200 35 Selling price per it Crrent demand per weok (unit Processing tme required 2,000 argin ratio 40P% 1.00 min. 0.50 min. Machine Al is the scarce resource because there is excess capacity on other machines Machine Al is being used at 100% of its capacity. Machine Al capacity is 2,400 minutes per week. Should Kaiser focus its efforts on Product 1 or 2 to maximize its contribution margin? Section5(20) The following information regarding Greenwich Company is available: Sales... $1,512,000 Operating expenses.... $252,000 Operating earningA Average invested capital Return on sales.D Capital turnover 40% Compute the iswers for items A-ID
Answer #1

Section 5(20) has been answered here, along with workings

Working Answer for 'A'
A Sales $          1,512,000
B Cost of Sales $              882,000
C Operating Expenses $              252,000
D = A - B - C Operating Earnings $              378,000
Working Answer for 'B'
A Sales $          1,512,000
B Capital Turnover 40%
C = A/B Average Invested Capital $          3,780,000
Working Answer for 'C'
A Operating Earnings $              378,000
B Average Invested Capital $          3,780,000
C = (A/B) x 100 ROI [Return on Investment] 10%
Working Answer for 'D'
A Operating Earnings $              378,000
B Sales $          1,512,000
C = (A/B) x 100 Return on Sales 25%